### v1

# Amortizing Perpetual Options

Abstract
In this work, we introduce amortizing perpetual options (AmPOs), a fungible variant of continuous-installment options suitable for exchange-based trading. Traditional installment options lapse when holders cease their payments, destroying fungibility across units of notional. AmPOs replace explicit installment payments and the need for lapsing logic with an implicit payment scheme via a deterministic decay in the claimable notional. This amortization ensures all units evolve identically, preserving fungibility. Under the Black-Scholes framework, AmPO valuation can be reduced to an equivalent vanilla perpetual American option on a dividend-paying asset. In this way, analytical expressions are possible for the exercise boundaries and risk-neutral valuations for calls and puts. These formulas and relations allow us to derive the Greeks and study comparative statics with respect to the amortization rate. Illustrative numerical case studies demonstrate how the amortization rate shapes option behavior and reveal the resulting tradeoffs in the effective volatility sensitivity.
1 Introduction
Perpetual contracts are experiencing a renaissance due to their adoption in decentralized finance and cryptocurrency markets. Though introduced in [17] for, e.g., a market on real estate indices, perpetual futures on cryptocurrencies are liquidly traded on a number of markets such as Hyperliquid and Binance. These contracts readily allow traders to purchase leveraged positions or to short assets which, otherwise, are technologically incapable of such a position; we refer the interested reader to, e.g., [1, 3]. However, perpetual futures only provide users with linear payoffs. Perpetual options, which replicate elements of the funding rate structure of perpetual futures, have recently been introduced in decentralized systems. We highlight everlasting options [19] and Panoptic [14] as specific perpetual option variants.
However, though perpetual futures have found wide usage in cryptocurrencies, perpetual options have not gained the same market share. Notably, the stochastic funding rates used in everlasting options and at Panoptic can be viewed as types of installment options. An installment option is one in which cash flows are exchanged between the long and short positions [8, 4]. Traditionally, these cash flows exist solely from the long to the short holders. Such options, however, only trade over-the-counter due to the nature of these contracts [4]. Specifically, installment options grant the holder a dual optionality: exercise or lapse. Because the lapsing logic comes from failure to make the installment payment in full, the notional units of these contracts are not identical and, therefore, cannot trade on an exchange. For the purposes of this work, we specifically highlight continuous-installment (CI) options which require a continuous stream of payments from the option holder to the option underwriter in order to keep the contract alive [12, 13].
The remainder of this paper is organized as follows. In Section 2, we introduce an exchange-tradable variant of CI options; this is accomplished by replacing the explicit installment payments with implicit amortization of the options notional. In Section 3, under the Black-Scholes framework, we study the risk-neutral valuation and Greeks by determining an equivalence with vanilla perpetual American options for a dividend-paying asset. In that same section, we also extend the Greeks to consider sensitivity of the option behavior to the contractual amortization rate. Section 4 then presents a brief discussion of applicability of this new option type for use in hedging for both traditional and decentralized finance settings. Section 5 concludes.
2 Amortizing Options
As highlighted within Section 1, CI options present a novel approach to perpetual contracts which enforce an explicit time-cost on the option holder [6, 12, 13]. For such options, the buyer pays an initial premium $V_{0}>0$ to the underwriter for a claimable notional of $N_{0}>0$ upon exercise. However, in contrast to traditional options, the holder also needs to pay continuous installments $c_{t}dt$ per unit of notional to the underwriter to keep the contract alive. In this way, the holder has two forms of optionality:
  1. 1.
    the right to exercise, and
  2. 2.
    the right to lapse by halting the installment payments.
Installment options have been offered over-the-counter in a few markets: “installment warrants on Australian stocks listed on the Australian Stock Exchange (ASX) [10, 4], a 10-year warrant with nine annual payments offered by Deutsche Bank [7], and so on,”  [12, Section 1]. However, because lapsing destroys fungibility across units of notional, i.e. partial lapsing by making a partial installment payment is not permissible, traditional CI options are not amenable to exchange-based trading.
Example 2.1.
Traditionally, as described in [12], CI options specify installments as either: (1) constant over time $c_{t}\equiv c>0$, or (2) a constant multiple of the underlying asset’s value $c_{t}\equiv\gamma S_{t}$ for underlying $S_{t}$ with $\gamma>0$.
Herein we propose a variant of CI options. Instead of directly paying the continuous installments, the option holder implicitly sells a fraction of his or her option holdings back to the underwriter at the current premium. Formally, if $N_{t}>0$ denotes the time $t$ claimable notional and $V_{t}>0$ denotes the option’s premium per unit notional, then funding the installment cost $c_{t}N_{t}dt$ with a sale of the notional leads to an exponential decay in the notional:
$dN_{t}=-q_{t}N_{t}dt,\quad q_{t}=c_{t}/V_{t}.$
Notably, because the notional is decreasing from these implicit payments, the requisite installment payments are also decreasing over time. In contrast to CI options in which the option underwriter is paid continuously in time, this option variant “pays” the underwriter via the decrease in his or her risk exposure.
Definition 2.2.
An amortizing perpetual option (AmPO) is an implicit payment, perpetual American CI option with installment costs $c_{t}=qV_{t}$ for some constant amortization rate $q>0$.
Remark 1.
While the fraction-of-premium installment payments ($c_{t}=qV_{t}$) implied by AmPO contracts is non-standard, it leads to a fixed, predictable amortization rate $q>0$ for the notional. In this way, the holder’s position deterministically decays over time, independent of the underlying market conditions. On the other hand, the underwriter is continuously compensated in proportion to the notional and fair-value exposure being carried; this holds because the implicit installment payments reduce that exposure in line with the contract’s value.
In contrast, under standard CI structures (as in Example 2.1), the amortization rate $c_{t}/V_{t}$ can spike or plummet with market conditions. For instance, if the underlying asset’s price remains constant, but its volatility rises, then $V_{t}$ increases while $c_{t}$ remains constant. Under such a scenario, the realized amortization rate would fall, leaving the underwriter with a disproportionately large remaining exposure.
Remark 2.
Alternative installment payments (e.g., as in Example 2.1) can be chosen instead, leading to time-varying amortization rates $c_{t}/V_{t}$. Regardless of the installment cost structure, amortizing CI variants become exchange tradable as every unit of notional is always identical and cannot lapse independently, i.e., all contract units are perfectly fungible. For such state-dependent amortization schedules, the valuation problem to be presented in Section 3.1 can be approached along the lines of [2]. For simplicity, herein we focus solely on the constant amortization case presented in Definition 2.2.
Remark 3.
Step options [15, 9] similarly feature payoffs with an effective notional that can decay over time. However, in contrast to the deterministic amortization of AmPOs, the decay in step options depends on the occupation time of the underlying, thus requiring path-dependent valuation and eliminating the fungibility required for exchange-based trading.
3 Risk-Neutral Valuation
Call Option Put Option
Price $\frac{K}{\alpha_{C}-1}\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}}$ $\frac{K}{1+\alpha_{P}}\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{\alpha_{P}}$
Delta $\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}-1}$ $\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{1+\alpha_{P}}$
Gamma $\frac{(\alpha_{C}-1)^{2}}{\alpha_{C}K}\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}-2}$ $\frac{\alpha_{P}K}{S_{0}^{2}}\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{\alpha_{P}}$
Theta 0 0
Vega $\frac{4C_{0}}{\sigma}\log\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)\left(\frac{(\alpha_{C}-2)r-q}{(2\alpha_{C}-1)\sigma^{2}+2r}\right)$ $\frac{4P_{0}}{\sigma}\log\left(\frac{(1+\alpha_{P})S_{0}}{\alpha_{P}K}\right)\left(\frac{(2+\alpha_{P})r+q}{(2\alpha_{P}+1)\sigma^{2}-2r}\right)$
Table 1: Summary table of AmPO risk-neutral valuation and Greeks for a constant notional exposure.
Within this section, we want to investigate the risk-neutral valuation of these AmPOs as presented in the prior section. To do so, we follow Assumption 3.1 throughout the remainder of this work.
Assumption 3.1.
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space satisfying the usual conditions. Consider a complete market with constant risk-free rate $r\geq 0$ and risky asset with price $S=(S_{t})_{t\geq 0}$. Let $\mathbb{Q}\sim\mathbb{P}$ denote the unique risk-neutral measure for this market and consider the $\mathbb{Q}$-Brownian motion (w.r.t. $(\mathcal{F}_{t})_{t\geq 0}$) $W=(W_{t})_{t\geq 0}$. For the remainder of this work, we will assume that $S$ follows a geometric Brownian motion under $\mathbb{Q}$:
$dS_{t}=S_{t}(rdt+\sigma dW_{t})$
with initial price $S_{0}>0$ and volatility $\sigma>0$.
3.1 Pricing
To determine the risk-neutral value of an AmPO, we first observe a fundamental symmetry in the value process. Holding an AmPO involves a decaying notional $N_{t}=e^{-qt}N_{0}$ on the options payoff $\Phi(\cdot)$. That is, from the holder’s perspective, the position’s value is $N_{t}V(S_{t})$ for AmPO option value $V(\cdot)$ with payoff $N_{\tau}\Phi(S_{\tau})$ at exercise time $\tau$. The amortization immediately shows up as an additional discounting term (i.e., an effective risk-free rate $r+q$); with this augmented risk-free rate, the amortization also functionally appears as a drag on the modified drift to the underlying asset (i.e., an effective dividend yield $q$). In this way, in Lemma 3.2, we demonstrate that AmPO valuation perfectly coincides with (vanilla) perpetual American options under modified parameters.
Lemma 3.2.
The AmPO with amortization rate $q>0$ is priced identically to the perpetual American option with the same payoff, effective risk-free rate $r+q$, and dividend yield $q$.
Proof.
Let $V:\mathbb{R}_{+}\to\mathbb{R}$ denote the price of one unit of the AmPO as a function of the underlying. Given an original notional position $N_{0}=1$ wlog, the AmPO contract’s value at time $t$ is given by $\Pi_{t}=e^{-qt}V(S_{t})$. In particular, the discounted value process $(e^{-rt}\Pi_{t})_{t\geq 0}$ must be a $\mathbb{Q}$-martingale. Applying Itô’s lemma, we recover the standard Black-Scholes differential equation:
$0=\frac{1}{2}\sigma^{2}S_{t}^{2}V^{\prime\prime}(S_{t})+rS_{t}V^{\prime}(S_{t})-(r+q)V(S_{t}).$
The desired result follows directly by comparing this ODE to the standard result for perpetual American options on dividend-paying assets (see, e.g., [11, Chapter 26.2]). ∎
Following Lemma 3.2, the pricing of call and put AmPOs trivially follow from the results of, e.g., [11, Chapter 26.2]. For this reason, we omit the proof for the pricing of call and put AmPOs.
Corollary 3.3 (Call Option Pricing).
Consider a call AmPO with strike $K>0$ and amortization rate $q>0$. The optimal exercise boundary is $\bar{S}_{C}=\frac{\alpha_{C}K}{\alpha_{C}-1}$ for $\alpha_{C}=\sqrt{\left(\frac{r}{\sigma^{2}}+\frac{1}{2}\right)^{2}+\frac{2(r+q)}{\sigma^{2}}}-\frac{r}{\sigma^{2}}+\frac{1}{2}>1$. Under optimal execution, the premium for this option is
$C_{0}=\frac{K}{\alpha_{C}-1}\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}}.$
Corollary 3.4 (Put Option Pricing).
Consider a put AmPO with strike $K>0$ and amortization rate $q>0$. The optimal exercise boundary is $\bar{S}_{P}=\frac{\alpha_{P}K}{1+\alpha_{P}}$ for $\alpha_{P}=\sqrt{\left(\frac{r}{\sigma^{2}}+\frac{1}{2}\right)^{2}+\frac{2(r+q)}{\sigma^{2}}}+\frac{r}{\sigma^{2}}-\frac{1}{2}>0$. Under optimal execution, the premium for this option is
$P_{0}=\frac{K}{1+\alpha_{P}}\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{\alpha_{P}}.$
We conclude this discussion of risk-neutral valuation by considering how the amortization $q$ can mimic an effective maturity date $T$. That is, for a given amortization rate, we can define the effective maturity as the time $T$ such that the option’s premium coincides with a standard, dated American option with the same underlying dynamics.
Example 3.5.
Consider an asset price following Assumption 3.1 with risk-free rate $r=5\%$, volatility $\sigma=50\%$, and initial price $S_{0}=100$. Consider the at-the-money call option (i.e., $K=100$) with amortization rate $q$. In Figure 1(a), we plot the maturity of the at-the-money (American) call option which has the same premium as the AmPO with amortization rate qq. [^1] As the assets under consideration here pay 0 dividends, American and European call options have the same premium. As expected, the effective maturity $T$ is decreasing in the amortization rate $q$. Notably, even though the effective maturity is decreasing, the effective notional ($e^{-qT}$) at the effective maturity is also decreasing in the amortization rate (see Figure 1(b)). For the selected parameters, even at high amortization rates ($q\approx 1$), the AmPO has over 65% of its original notional value remaining at the time that the equivalent dated option would expire.
(a) The effective maturity date $T$ as a function of the amortization rate $q$.
(b) The effective notional ($e^{-qT}$) of the AmPO with amortization rate $q$ at its effective maturity date $T$.
Figure 1: Example 3.5: Comparison of AmPO and dated at-the-money call options.
3.2 Greeks
As call and put AmPOs can be valued via an equivalence relation with perpetual American options as provided in Section 3.1, we can follow standard results to determine the dependence of these structures on the underlying system parameters. The main results from this section are summarized in Table 1. Throughout this section we omit the proofs as they follow directly from the aforementioned relation to vanilla perpetual American options.
Remark 4.
Herein we consider the Greeks of the valuation for a constant notional exposure, i.e., dependence of the explicit premia $C_{0}$ and $P_{0}$ for calls and puts, respectively. However, we note that the realized exposure is amortizing for an investor that is either long or short AmPOs. Therefore, in practice, the realized Greeks should be discounted by the amortization factor appropriately. (Further discussions will be provided for Theta in Remark 5 below.)
Corollary 3.6 (Delta).
Consider call and put AmPOs with strike $K>0$ and amortization rate $q>0$. The Delta for call and put options are provided, respectively, by:
$\Delta_{C}:=\frac{\partial C_{0}}{\partial S_{0}}=\begin{cases}\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}-1}&\text{if }S_{0}\leq\bar{S}_{C}\\ 1&\text{if }S_{0}>\bar{S}_{C}\end{cases}\quad\text{ and }\quad\Delta_{P}:=\frac{\partial P_{0}}{\partial S_{0}}=\begin{cases}-\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{1+\alpha_{P}}&\text{if }S_{0}\geq\bar{S}_{P}\\ -1&\text{if }S_{0}<\bar{S}_{P}.\end{cases}$
Furthermore, $\Delta_{C}>0$ and $\Delta_{P}<0$ for any $S_{0}>0$.
Corollary 3.7 (Gamma).
Consider call and put AmPOs with strike $K>0$ and amortization rate $q>0$. The Gamma for call and put options are provided, respectively, by:
$\Gamma_{C}:=\frac{\partial^{2}C_{0}}{\partial S_{0}^{2}}=\begin{cases}\frac{(\alpha_{C}-1)^{2}}{\alpha_{C}K}\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}-2}&\text{if }S_{0}\leq\bar{S}_{C}\\ 0&\text{if }S_{0}>\bar{S}_{C}\end{cases}\;\text{ and }\;\Gamma_{P}:=\frac{\partial^{2}P_{0}}{\partial S_{0}^{2}}=\begin{cases}\frac{\alpha_{P}K}{S_{0}^{2}}\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{\alpha_{P}}&\text{if }S_{0}\geq\bar{S}_{P}\\ 0&\text{if }S_{0}<\bar{S}_{P}.\end{cases}$
Furthermore, $\Gamma_{C},\Gamma_{P}\geq 0$ for any $S_{0}>0$.
Remark 5 (Theta).
Theta, typically, refers to the dependence of the premium on the “time to maturity”. Because AmPOs are perpetual options, the risk-neutral price exhibits zero explicit time decay (e.g., $\frac{\partial C_{0}}{\partial t}=0$ for call options). However, due to the amortization, the holder or underwriter of such options will, in fact, have an economic exposure that depends on time. This decay is directly computable as $qC_{0}<0$ or $qP_{0}<0$ for call and put options, respectively. It is exactly this non-zero Theta(-like) decay that distinguishes AmPOs from vanilla perpetual options and creates an effective realized time exposure.
Corollary 3.8 (Vega).
Consider call and put AmPOs with strike $K>0$ and amortization rate $q>0$. The Vega for call and put options are provided, respectively, by:
$\nu_{C}$ $=\frac{\partial C_{0}}{\partial\sigma}=\begin{cases}\frac{4C_{0}}{\sigma}\log\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)\left(\frac{(\alpha_{C}-2)r-q}{(2\alpha_{C}-1)\sigma^{2}+2r}\right)&\text{if }S_{0}\leq\bar{S}_{C}\\ 0&\text{if }S_{0}>\bar{S}_{C}\end{cases}$
$\nu_{P}$ $=\frac{\partial P_{0}}{\partial\sigma}=\begin{cases}\frac{4P_{0}}{\sigma}\log\left(\frac{(1+\alpha_{P})S_{0}}{\alpha_{P}K}\right)\left(\frac{(2+\alpha_{P})r+q}{(2\alpha_{P}+1)\sigma^{2}-2r}\right)&\text{if }S_{0}\geq\bar{S}_{P}\\ 0&\text{if }S_{0}<\bar{S}_{P}.\end{cases}$
Furthermore, $\nu_{C},\nu_{P}\geq 0$ for any $S_{0}>0$.
Example 3.9.
Consider the same setting as Example 3.5. Again comparing AmPOs to their effective dated counterparts, we want to consider the relative safety of AmPOs. In Figure 2(a), we plot the “safety ratio” between the AmPO Gamma and the Gamma of the effectively dated (American) call option. The lower this ratio, the more stable the option price is to fluctuations in the underlying. Notably, this ratio is increasing in the amortization rate, but stays below 80% for even extremely high amortization rates (e.g., $q=1$). Similarly, in Figure 2(b), we consider the “cost efficiency” of AmPOs by computing the ratio between the economic Theta (i.e., $qC_{0}$) from amortization against the Theta of the effectively dated option. Here, again, we find that the dated options lose value faster even at high amortization rates with a maximum observed efficiency ratio (at $q=1$) of under 75%.
(a) The ratio of the AmPO Gamma against its effectively dated peer.
(b) The ratio of the AmPO (economic) Theta against its effectively dated peer.
Figure 2: Example 3.9: Comparison of AmPO and dated at-the-money call options.
3.3 Comparative Statistics on Amortization Rate
Finally, as the amortization rate $q$ appears directly within both the effective risk free rate ($r+q$) and effective dividend yield ($q$) of the equivalently valued perpetual American option, the choice of this contractual parameter can have a strong influence on the valuation and Greeks of AmPOs. However, since we have proposed AmPOs to maintain a constant $q$ throughout the lifetime of the option, we do not want to consider this dependence in the same way as the Greeks (i.e., for hedging purposes), but rather as a way to understand the implications and tradeoffs of choosing the $q$ for a contract. In doing so, we will focus on three aspects of the AmPO contracts: the risk-neutral valuation; the optimal exercise boundary; and the Vega. Throughout this section we set $\bar{\alpha}:=(\alpha_{C}+\alpha_{P})/2>0$ for notational simplicity.
Assumption 3.10.
For simplicity of notation, throughout this section we consider $S_{0}\leq\bar{S}_{C}$ for call options and $S_{0}\geq\bar{S}_{P}$ for put options.
First, as expected, increasing $q$ accelerates economic amortization, reducing the value for the holder and, therefore, also the premium. This is provided in Corollary 3.11 below.
Corollary 3.11.
Consider call and put AmPOs with strike $K>0$. The premium for both option types is decreasing in the amortization rate $q>0$:
$\frac{\partial C_{0}}{\partial q}=\frac{C_{0}}{\sigma^{2}\bar{\alpha}}\log\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)\leq 0\quad\text{ and }\quad\frac{\partial P_{0}}{\partial q}=-\frac{P_{0}}{\sigma^{2}\bar{\alpha}}\log\left(\frac{(1+\alpha_{P})S_{0}}{\alpha_{P}K}\right)\leq 0.$
Proof.
Consider the call AmPO; we omit the proof of put AmPOs as the arguments follow comparably. Recall from Corollary 3.3 that the premium of the call AmPO $C_{0}$ depends on the amortization rate $q$ solely through $\alpha_{C}$. Thus, through an application of the chain rule, we find $\frac{\partial C_{0}}{\partial q}=\frac{\partial C_{0}}{\partial\alpha_{C}}\frac{\partial\alpha_{C}}{\partial q}$. By differentiation, we find
$\frac{\partial C_{0}}{\partial\alpha_{C}}=C_{0}\log\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)\leq 0\quad\text{ and }\quad\frac{\partial\alpha_{C}}{\partial q}=\frac{1}{\sigma^{2}\bar{\alpha}}>0$
thus completing the proof. ∎
Remark 6.
Within Corollary 3.11, we found that the premium for both AmPO calls and puts are decreasing in the amortization rate $q$. In fact, it can easily be seen that the limiting behavior of these options is to vanilla perpetual American options as $q\searrow 0$ and to the intrinsic value of the option as $q\nearrow\infty$.
As the option premium decreases in the amortization rate, the optimal exercise boundary is similarly decaying towards the strike price (i.e., decreasing for call options and increasing for put options). In this way, the optimal exercise becomes more conservative as the amortization accelerates.
Corollary 3.12.
Consider call and put AmPOs with strike $K>0$. The optimal exercise boundary is decreasing to $K$ for call options and increasing to $K$ for put options as $q$ increases, i.e.,
$\frac{\partial\bar{S}_{C}}{\partial q}=-\frac{K}{\sigma^{2}(\alpha_{C}-1)^{2}\bar{\alpha}}\leq 0\quad\text{ and }\quad\frac{\partial\bar{S}_{P}}{\partial q}=\frac{K}{\sigma^{2}(1+\alpha_{P})^{2}\bar{\alpha}}\geq 0$
with $\lim_{q\nearrow\infty}\bar{S}_{C}=K$ and $\lim_{q\nearrow\infty}\bar{S}_{P}=K$.
Proof.
As with the proof of Corollary 3.11, this result trivially follows from an application of the chain rule as $\bar{S}_{C},\bar{S}_{P}$ depend only on $q$ through $\alpha_{C},\alpha_{P}$ respectively. ∎
Finally, we find that not only does the AmPO premium decrease and the optimal exercise boundary decay, as amortization accelerates, but we also see that the Vega of these options also drops. In this way, higher amortization rates can lead to option prices that are less sensitive to the underlying market dynamics. In fact, as highlighted in Remark 6, as the amortization rate approaches $\infty$, the premium of these options approaches the intrinsic value which is insensitive to the volatility.
Corollary 3.13.
Consider call and put AmPOs with strike $K>0$. The Vega is decreasing in the amortization rate $q$.
Proof.
Fix the amortization rate $q>0$.
  • Call Options: Consider the dependence of $C_{0}$ on $q$ as provided in Corollary 3.11. Notably, the dependence on $\sigma$ occurs only through $\alpha_{C}$ and $\bar{\alpha}$. Thus, from an application of the chain rule, we find
    $\frac{\partial^{2}C_{0}}{\partial\sigma\partial q}=\left(\frac{\partial}{\partial\alpha_{C}}\frac{\partial C_{0}}{\partial q}\right)\frac{\partial\alpha_{C}}{\partial\sigma}+\left(\frac{\partial}{\partial\bar{\alpha}}\frac{\partial C_{0}}{\partial q}\right)\frac{\partial\bar{\alpha}}{\partial\sigma}.$
    Expanding out each individual term we conclude
    $\frac{\partial}{\partial\alpha_{C}}\frac{\partial C_{0}}{\partial q}$ $=\frac{C_{0}}{\sigma^{2}\alpha_{C}\bar{\alpha}}\left[\alpha_{C}(\alpha_{C}-1)\log\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{2}+1\right]>0$
    $\frac{\partial\alpha_{C}}{\partial\sigma}$ $=\frac{1}{\sigma^{3}}\left[2r-\frac{2r^{2}+\sigma^{2}(3r+2q)}{\sigma^{2}\bar{\alpha}}\right]$
    $\frac{\partial}{\partial\bar{\alpha}}\frac{\partial C_{0}}{\partial q}$ $=-\frac{C_{0}}{\sigma^{2}\bar{\alpha}^{2}}\log\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)>0$
    $\frac{\partial\bar{\alpha}}{\partial\sigma}$ $=-\frac{2r^{2}+\sigma^{2}(3r+2q)}{\sigma^{5}\bar{\alpha}}<0.$
    Thus the stated monotonicity of Vega holds if $\frac{\partial\alpha_{C}}{\partial\sigma}<0$. In fact, this is easily verified by observing $4r^{2}\sigma^{4}\bar{\alpha}^{2}<(2r^{2}+\sigma^{2}(3r+2q))^{2}$.
  • Put Options: Consider the dependence of $P_{0}$ on $q$ as provided in Corollary 3.11. Notably, the dependence on $\sigma$ occurs only through $\alpha_{P}$ and $\bar{\alpha}$. Thus, from an application of the chain rule, we find
    $\frac{\partial^{2}P_{0}}{\partial\sigma\partial q}=\left(\frac{\partial}{\partial\alpha_{P}}\frac{\partial P_{0}}{\partial q}\right)\frac{\partial\alpha_{P}}{\partial\sigma}+\left(\frac{\partial}{\partial\bar{\alpha}}\frac{\partial P_{0}}{\partial q}\right)\frac{\partial\bar{\alpha}}{\partial\sigma}.$
    Expanding out each individual term we immediately reach the desired monotonicity:
    $\frac{\partial}{\partial\alpha_{P}}\frac{\partial P_{0}}{\partial q}$ $=\frac{P_{0}}{\sigma^{2}\bar{\alpha}}\left[\log\left(\frac{(1+\alpha_{P})S_{0}}{\alpha_{P}K}\right)^{2}+\frac{1}{(1+\alpha_{P})\alpha_{P}}\right]>0$
    $\frac{\partial\alpha_{P}}{\partial\sigma}$ $=-\frac{1}{\sigma^{3}}\left[2r+\frac{2r^{2}+\sigma^{2}(3r+2q)}{\sigma^{2}\bar{\alpha}}\right]<0$
    $\frac{\partial}{\partial\bar{\alpha}}\frac{\partial P_{0}}{\partial q}$ $=\frac{P_{0}}{\sigma^{2}\bar{\alpha}^{2}}\log\left(\frac{(1+\alpha_{P})S_{0}}{\alpha_{P}K}\right)>0$
    $\frac{\partial\bar{\alpha}}{\partial\sigma}$ $=-\frac{2r^{2}+\sigma^{2}(3r+2q)}{\sigma^{5}\bar{\alpha}}<0.$
In the prior Examples 3.5 and 3.9, we have analyzed how AmPOs compare with dated options. In the following example, we wish to consider how selecting the amortization rate can be used to optimize a volatility-based strategy.
Example 3.14.
Consider the setting of Example 3.5. Recall from Corollary 3.11 and Corollary 3.13 that both the premium and Vega of an AmPO are decreasing in the amortization rate $q$. As the realized Vega of a position depends on the budget actually invested, we now want to consider how the ratio of the Vega to premium behaves to consider a simplified optimization problem: consider an investor who wants to take as leveraged a volatility position as possible using AmPOs with a \$100 budget constraint. To do so, we consider three strategies: bullish volatility with call options (Figure 3(a)); bearish volatility with put options (Figure 3(b)); and neutral volatility with an at-the-money straddle (Figure 3(c)). Notably, though bullish and neutral strategies have increasing positional Vegas, the bearish strategy has a non-monotonic behavior coming from the simultaneous drop in both premium and Vega. For the current parameter setting, the optimal bearish volatility strategy (which also provides the maximum positional Vega among all 3 strategies) is with an amortization rate of $q^{*}\approx 14.26\%$.
(a) Call options
(b) Put options
(c) Straddle
Figure 3: Example 3.14: Vega of a \$100 position in call, put, and straddle strategies.
4 Discussion and Applications
Thus far, we have considered the construction and valuation of AmPOs. Within this section, we will briefly consider use cases for these novel perpetual options. In doing so, we will consider applications both within traditional financial markets and decentralized, i.e., blockchain-based, markets.
4.1 Traditional Finance
Consider that installment options have been traded in over-the-counter markets already [12]. However, as noted in Section 2, such options are non-fungible because of the specific requirements of installment payments and lapsing logic. In contrast, AmPOs provide an option type that can be transacted on a central limit order book. This exchange-based tradability permits more robust price-discovery and the possibility of more frequent transactions (due to, e.g., the ability to close long and short positions easily with market operations).
Notably, due to the deterministic cost-of-carry of AmPOs, large institutional investors (e.g., pension plans) can hold a large perpetual protective position while only needing to pay small re-upping charges to maintain the desired notional exposure. This is in contrast to dated options which require rolls around the expiry date. Such rolling operations become high-risk events due to the timing involved as, e.g., the Greeks of at-the-money options can explode near expiry.
4.2 Decentralized Finance
While exchange-tradability is beneficial in traditional markets, it is essential for use in decentralized finance (DeFi) due to the unique nature of blockchain technologies. Though non-fungible tokens exist in decentralized finance, they are generally traded in illiquid markets, such as OpenSea, because every contract would need to find a specific counterparty. In comparison, fungible tokens trade in liquid off-chain limit order books or on-chain automated market makers.
Automated market makers function by holding reserves of two or more assets against which anyone can transact [5]. These asset reserves are supplied by liquidity providers in exchange for the fees earned on every transaction. However, though the liquidity providers earn fees, their supplied assets are at risk of being sold to liquidity takers. Because of adverse selection effects, the effective position held by the liquidity providers (less the fees) is strictly worse than if replicated in an external market; this economic difference has been termed the loss-versus-rebalancing [16]. Notably, as a variant of CI options, AmPOs would be readily usable to hedge against this loss-versus-rebalancing [18] without requiring rolling-based risks.
5 Conclusion
Within this work, we introduced amortizing perpetual options, a fungible variant of continuous-installment options which are, therefore, suitable for exchange-based trading. Under a constant amortization rate, we found an equivalence relation between AmPOs and vanilla perpetual American options on dividend-paying assets. This relation permits standard analytical valuations and, thus, allowed us to characterize the Greeks and analyze the sensitivities of option behavior to the contractual amortization rate.
Our analysis in Section 3 assumed the complete market structure of the Black-Scholes framework with the constant amortization rate. Extending the study of AmPOs to incomplete markets and stochastic amortization rates presents several natural directions for future work. For instance, one could investigate the optimal amortization rate for long and short positions as there is a natural tradeoff between a higher amortization rate and lower initial premium (Corollary 3.11). In this way, an equilibrium amortization rate can arise from utility-indifference or heterogeneous-agent considerations. Such extensions would provide a framework for designing amortization rates in practice.